OpenStudy (anonymous):

Related rates problem... A tumor the shape of a sphere is found to have its diameter increasing at a rate of 0.2 cm per month. The mass, in grams, of the tumor is measured as m = (8/15)d 3, where d is the diameter of the tumor. How fast is the mass of the tumor increasing (in grams per month) when the diameter measures 2.4 cm?

OpenStudy (jonnyvonny):

Is the problem m=(8/15)*3d?

OpenStudy (jonnyvonny):

equation*, rather

OpenStudy (jonnyvonny):

Or \[m=(8/15)*d^3?\]

OpenStudy (anonymous):

Oh, sorry! It's supposed to be (8/15)d^3

OpenStudy (anonymous):

m = (8/15)*d^3

OpenStudy (jonnyvonny):

What do we need to find?

OpenStudy (jonnyvonny):

We know d, and dd/dt.

OpenStudy (anonymous):

Are you going to walk me through the steps? :) It asks to find the time, g/month, of the tumor's rate of growth. All I've got so far is... m=(8/15)*d^3 = (8/15)*(2.4cm)^3 = 7.3728cm^3...or 7.3728g...What do I need to do next?...

OpenStudy (jonnyvonny):

Well, anyways, I g2g.. so what you do is take the derivative of the function, as is and get \[dm/dt=(8/15)*3*d^2*dd/dt\] From then, you just plug in the values they provide, and get the answer

OpenStudy (anonymous):

Oh I see! Thanks so much!! :)

OpenStudy (jonnyvonny):

Yw, I'm back, do you seek for more elaboration for this problem? Sorry for sounding rude; I had to leave.

OpenStudy (anonymous):

No no, I didn't mind at all :) I figured it out with your help. Thanks!